### v1

@@ 3.2 Systematic uncertainties of event generation in DIS
The first source encompasses several perturbative ingredients of the calculation, i.e. the evaluation of hard matrix elements at various multiplicities in merged samples, as well as parton shower evolution. We follow the established approach of estimating the size of missing higher perturbative orders by examining the residual dependence of resulting observables on the renormalization and factorization scales. We note though that this approach has severe limitations and especially in the context of even higher order calculations it might be appropriate to investigate other sources of uncertainty [63, 64, 65, 66, 67]. Here, we perform 7-point variations around a central choice $(\mu_{F,0},\mu_{R,0})$, considering all ratios
@@
We apply these variations in the matrix-element calculation and the parton-shower evolution simultaneously. Ultimately, we build our estimate of the perturbative uncertainty from the envelope of all scale choices. It is worth stressing that, in this procedure, we are varying scales appearing in the parton shower, which affects the evaluation of $\alpha_{\text{s}}$ and the PDF ratios. Providing more extensive parton shower uncertainties is an open topic of research, complicated by the fact that parton showers include significant subleading contributions [68]. We here stay restricted to the traditional scale variations. S HERPA performs this operation on-the-fly [69], which significantly reduces computing costs compared to running separate simulations for each scale setting.
@@
The second source of uncertainty relates to the non-perturbative modeling of parton fragmentation into hadrons. These uncertainties arise from phenomenological models rather than first-principle calculations, leading to a significant model dependence. To evaluate fragmentation corrections, we use S HERPA ’s internal cluster hadronization model [56] and the Lund String model via an interface to P YTHIA 8 [57]. The symmetrized difference between these two models is taken as the uncertaintyNote this procedure can in principle result in uncertainties covering negative (differential) cross-sections. These are clearly unphysical and we cut off the envelope at zero., using the cluster model as the central value.
@@
Another aspect of non-perturbative modeling involves tuning model parameters to data. We use the replica tunes from Ref. [70] to estimate uncertainties related to the tuning of final state fragmentation parameters and the variations from Ref. [71] for parameters governing cluster decays involving beam remnants. This introduces two sets of variations within the cluster model itself, for which we take the respective envelopes for each histogram as the related uncertainty. As a result, three sources of non-perturbative uncertainty arise: model choice, final state fragmentation tune, and beam fragmentation tune. Arguably the different sources of uncertainty should be combined. We however find that usually only one, in most cases the model choice, dominates the overall uncertainty. We hence only show the three sources independently, without a combination.
@@
Since estimating non-perturbative uncertainties requires independent Monte Carlo runs for each parameter set, and for both models, performing this study for the full MEPS@NLO, MEPS@LO and MC@NLO setups is infeasible. Instead, we derive the relative uncertainty based on parton showered leading-order matrix elements merged together in the MEPS@LO approach. A merged approach is still beneficial to ensure sufficient phase-space coverage. In the following we only treat the results of two tunes or models as different if we can statistically distinguish them at a level of 3 standard deviations, and otherwise indicate the statistical error that serves as an upper bound on the uncertainty.
@@
We make no attempt to directly apply this uncertainty to higher- or lower-precision simulations, though one could envision a procedure based on transfer matrices introduced in Ref. [72]. Additionally, we do not combine non-perturbative and perturbative uncertainty estimates but rather focus on comparing their relative sizes in the following analysis.
@@ 4 Event Generation for HERA and MEPS@NLO in charged current DIS
Figure 2: Differential distributions for $Q^{2}$ and $x$ at LO, MC@NLO, MEPS@LO and MEPS@NLO in inclusive CC DIS at the HERA. The lower panels show the ratio between data and the merged MEPS@LO and MEPS@NLO predictions, while we omit the ratio for the LO and MC@NLO simulations for clarity. Experimental data points are from Ref. [73].
The S HERPA framework has been extensively used in the past to describe neutral current DIS at HERA, see for example Ref. [12]. Early versions of S HERPA 3 were also used in a variety of recent event shape and jet analyses by the H1 collaboration[74, 75, 76, 46, 77], usually finding satisfactory agreement with the data. A MEPS@NLO setup similar to the one used here was compared to resummed calculations for event shapes in Refs. [71, 78]. The latter also determined part of the current tune of non-perturbative parameters. With this, we consider the framework and MEPS@NLO to be sufficiently validated in neutral current DIS, and we focus in the following on the first detailed comparison to CC data from HERA.
[…]
A caveat for CC DIS is that, in contrast to NC DIS, it is impossible to use the final state neutrino to reconstruct $q$. Therefore, the measurements of Eqs. (2) involve the reconstruction of the full hadronic final state under the condition of large missing $p_{t}$ from the invisible high-energy neutrino. The details of this procedure are discussed in Refs. [73, 79]. At the hadronic simulation level, we ignore this complication and consider the recoil neutrino as known for the determination of $Q^{2}$, $x$ and $y$. In the analysis, events are selected following the experimental phase space cuts:
@@
Figure 3: Differential distributions for $\eta_{\mathrm{jet}}$ and $E_{T,\mathrm{jet}}$ at LO, MC@NLO, MEPS@LO and MEPS@NLO in inclusive CC DIS at the HERA. The lower panels show the ratio between data and the merged MEPS@LO and MEPS@NLO predictions, while we omit the ratio for the LO and MC@NLO simulations for clarity. Experimental data points are from Ref. [73].
@@
Figure 4: Multijet distributions differential in the mass of the two (left) or three (right) leading jets at LO, MC@NLO, MEPS@LO and MEPS@NLO in inclusive CC DIS at the HERA. The lower panels show the ratio between data and the merged MEPS@LO and MEPS@NLO predictions, while we omit the ratio for the LO and MC@NLO simulations for clarity. Experimental data points are from Ref. [73].
In Figs. 2 and 3, we reproduce the differential cross-sections for inclusive-jet production as functions of $Q^{2}$, $x$, and leading-jet $\eta$ and $E_{\perp}$. Due to the relatively large $Q^{2}$ requirement, higher-order corrections are rather small, and we find the simulations at all accuracy levels to be consistent with each other and with the data. In particular, the lack of a sizable shift between the LO and MEPS@NLO predictions can be attributed to the high $Q^{2}$ cut. This requirement restricts the phase space to regions where merging effects remain small. The lower panels, we show the ratio of the merged simulations to data. Since the central values for all Monte Carlo accuracies are very consistent, we do not show ratios for the LO and MC@NLO runs to increase legibility. For the two remaining runs, we also plot the band obtained from scale variations. Notably, while the central value is almost unchanged when going from MEPS@LO to MEPS@NLO, the scale dependence is reduced dramatically. This leads to some visible systematic trend of the MEPS@NLO simulation against data, especially in the transverse energy distribution, but the data is not precise enough there for a definite conclusion.
In Fig. 4, the dijet and three-jet invariant mass cross-sections are presented. The different Monte Carlo samples are again very consistent with each other. For the dijet mass, some deviations appear for very small and very large masses, although only those at small masses have any statistical significance. In the range $20\,\text{GeV}<M_{jj}<120\,\text{GeV}$, the Monte Carlo predictions are consistent with the data. In the three-jet mass case, prediction and data are consistent over the full measured range, albeit admittedly with significantly larger uncertainties on the data. The scale variations are reduced in both cases, although they remain sizable in the three-jet case.
@@ 5 Event generation for the EIC and its systematic uncertainties
Furthermore, we compare jet properties obtained using the $k_{t}$, anti-$k_{t}$, and C ENTAURO clustering algorithms, highlighting differences in jet multiplicities and leading jet transverse momentum. Finally, to probe the hadronic final state, we study the 1-jettiness distribution, presented in slices of $Q^{2}$ and $x$.
@@
[…]
@@ 5.1 Neutral Current
Figure 5: Differential distributions for $Q^{2}$ (top left), $x$ (top right), $y$ (bottom left) and lepton transverse momentum $p_{\text{T}}^{\mathrm{lep}}$ (bottom right) in NC DIS at the EIC, comparing LO+PS, MC@NLO, MEPS@LO and MEPS@NLO predictions. The main panels show results with scale variation uncertainties for all but the LO+PS calculation. The first ratio panel displays the relative deviation from MEPS@NLO to highlight merging effects. The second ratio shows the relative scale uncertainties.
We consider the neutral current setup first. In this case, we place a small virtuality cut, $Q^{2}>6\,\text{GeV}^{2}$. We present the global DIS observables $Q^{2}$, $x$, $y$ and $p_{\text{T}}^{\mathrm{lep}}$ in Fig. 5. For $Q^{2}$ and $x$, the merged calculations exhibit distinct behavior from the matched MC@NLO approach. In the $Q^{2}$ distribution (top left), we observe that for $Q^{2}<20\,\text{GeV}^{2}$, the MEPS@NLO prediction exceeds LO and MC@NLO of up to 30%, while showing a more modest 10% enhancement over MEPS@LO. This enhancement exceeds the scale uncertainty bands between MEPS@NLO and MC@NLO but remains within the MEPS@LO uncertainties. The predictions converge in the intermediate region $20\,\text{GeV}^{2}\leq Q^{2}\leq 5\cdot 10^{3}\,\text{GeV}^{2}$. At higher virtualities ($Q^{2}>5\cdot 10^{3}\,\text{GeV}^{2}$), we observe a systematic 10% deficit in the LO and MEPS@LO predictions relative to the NLO-accurate calculations, showing the importance of the virtual $\mathcal{O}(\alpha_{\text{s}})$ corrections in this regime.
Similar patterns can be seen in the $x$ and $p_{\text{T}}^{\mathrm{lep}}$ distributions (top and bottom right). Multi-jet merging effects dominate at $x<1.5\cdot 10^{-3}$ ($p_{\text{T}}^{\mathrm{lep}}<5\,\text{GeV}$), while NLO matching becomes crucial at $x>0.7$ ($p_{\text{T}}^{\mathrm{lep}}>65\,\text{GeV}$). The MEPS@NLO calculation shows reduced scale uncertainties compared to MEPS@LO in the low-$x$ and low-$p_{\text{T}}^{\mathrm{lep}}$ regions. It also shows smaller uncertainties than MC@NLO in the intermediate ranges of around $Q^{2}\approx 10^{2}\,\text{GeV}^{2}$ and $x\approx 10^{-2}$ as a consequence of calculating the $2\to 3$ process to NLO. The inelasticity distribution (bottom left) reveals good agreement between MEPS@NLO and MEPS@LO across the full range, while LO and MC@NLO predictions are uniformly suppressed by approximately 25%, though this deviation remains within the combined scale uncertainty bands.
@@
Figure 6: Jet multiplicity $N_{\mathrm{jet}}$ (top row) and leading jet transverse momentum $p_{T}^{\mathrm{jet}}$ (bottom row) distributions for anti-$k_{t}$ (left), $k_{t}$ (middle) and C ENTAURO (right) jet algorithms in NC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from non-perturbative model and tuning variations. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@
Figure 6 presents the jet multiplicity and leading jet transverse momentum distributions for three different clustering algorithms. While MEPS@NLO and MEPS@LO predictions show good concordance, both MC@NLO and LO calculations yield systematically lower rates. Despite reduced scale uncertainties in MEPS@NLO compared to MEPS@LO, both merged calculations exhibit significant scale variations, reflecting the LO accuracy of higher-multiplicity configurations that dominate the low-$Q^{2}$ regime and that contribute significantly in these observables. It is also worth noting that the C ENTAURO clustering algorithm leads to smaller higher-order corrections in this analysis. Indeed, the LO simulation is unable to fill the whole phase space. For the $k_{t}$ and anti-$k_{t}$, the MC@NLO sample is still off, compared to the merged ones, by a $K$-factor of about 1.5 for the low-$p_{T}$ region. Conversely, the C ENTAURO algorithm shows much smaller $K$-factors, in fact only a factor of 1.2 for the low-$p_{T}$ region with respect to the MEPS@NLO prediction.
@@
Hadronization uncertainties are generally smaller than the perturbative uncertainties and, as one might have expected, are largest for small $p_{T}$ (up to $5\%$) and for large jet multiplicities $(1-2\%)$. In the latter, we find that only a variation of the fragmentation model induces a significant effect, while replica tunes within the cluster model are negligible. Instead, in the $p_{T}$ spectrum, fragmentation model uncertainty is still the largest non-perturbative contribution for the anti-$k_{t}$ and $k_{t}$ jets, while variations from replica tunes are larger than model ones for C ENTAURO jets, although both are especially small. Beyond small values of $p_{T}$, the difference between hadronization models becomes unresolvable within the accumulated statistics. Instead, the combined statistical uncertainty from the two models is taken as an overestimate. These values are still comparable or subdominant to the perturbative uncertainties in the same part of the spectrum.
@@
Figure 7: Differential distributions of 1-jettiness $\tau$ in different bins of $Q^{2}$ and $x$ in NC DIS at the EIC. The total MEPS@NLO cross-section for each $x$-$Q^{2}$ bin is reported on top of the corresponding plot.
@@
We finally consider 1-jettiness, an event shape variable that is sensitive to the underlying parton dynamics and can be used, e.g., for global fits of the strong coupling constant $\alpha_{\text{s}}$ [81]. In Ref. [82], predictions of 1-jettiness at the EIC were computed with NNLO+N3LL accuracy. Here, we present in Figure 7 an overview of the 1-jettiness $\tau$ across bins of $x$ (left to right) and $Q^{2}$ (top to bottom). For the sake of presentation we omit the ratio panels here, the full plots can be found in the appendix. Overall, the shape of the distributions agree well between the different configurations. Similar to the behavior at the recent H1 measurement [76], we observe that the high-$Q^{2}$ has a pronounced peak at small $\tau$ values, while the low-$Q^{2}$ distributions tend to higher values, which underlines the importance of high-multiplicity matrix elements in the latter regime.
@@
For a more thorough discussion, we show the full breakdown of ratios and uncertainties in Fig. 8 for three different $x$ regions in the $31.610<Q^{2}/\,\text{GeV}^{2}<$ 31.6 range. Plots for the other $x$-$Q^{2}$ ranges in Fig. 7 can be found in App. A.1. We observe that the non-perturbative uncertainties are non-negligible for small $\tau$ values and of similar size as the MC@NLO and MEPS@NLO scale uncertainties in that region. The same qualitative behavior is visible in both the tuning variations and the difference between string and cluster models, while the latter are significantly larger. This is the case for both tuning uncertainties and variations of the hadronization model, where the latter are dominating. For larger values of 1-jettiness, either of the considered non-perturbative uncertainties are negligible, consistent with the observations in [71]. Furthermore, the non-perturbative uncertainties are—unsurprisingly—largest for small values of $Q^{2}$ and $x$. The shape of the distributions at small $Q^{2}$ underlines the importance of multi-parton topologies in this region.

### v2

@@ 3.2 Systematic uncertainties of event generation in DIS
The first source encompasses several perturbative ingredients of the calculation, i.e. the evaluation of hard matrix elements at various multiplicities in merged samples, as well as parton shower evolution. We follow the established approach of estimating the size of missing higher perturbative orders by examining the residual dependence of resulting observables on the renormalization and factorization scales. We note, though, that this approach has severe limitations and especially in the context of even higher-order calculations it might be appropriate to investigate other sources of uncertainty [70, 71, 72, 73, 74]. Here, we perform seven-point variations around a central choice $(\mu_{F,0},\mu_{R,0})$, considering all ratios
@@
We apply these variations in the matrix-element calculation and the parton-shower evolution simultaneously. Ultimately, we build our estimate of the perturbative uncertainty from the envelope of all scale choices. It is worth stressing that, in this procedure, we are varying scales appearing in the parton shower, which affects the evaluation of $\alpha_{\text{s}}$ and the PDF ratios. Providing more extensive parton shower uncertainties is an open topic of research, complicated by the fact that parton showers include significant subleading contributions [32]. We stay restricted here to the traditional scale variations.
The particular choice of the merging scale, while arbitrary in principle, also affects the perturbative side of a merged simulation. In our approach, see Eq. (9), it is dynamically determined by the two parameters $\bar{Q}_{\text{cut}}$ and $S_{\text{DIS}}$. To asses the related uncertainty, we only vary the resulting scale $Q_{\text{cut}}$ here by a factors $1/2$ and $2$ and form an envelope of those predictions. Note this procedure is equivalent to varying only $\bar{Q}_{\text{cut}}$ at large $Q^{2}$, where $\bar{Q}_{\text{cut}}$ effectively acts as a fixed scale, or only $S_{\text{DIS}}$ at small $Q^{2}$ where the effective scale is just proportional to that.
S HERPA performs this operation on the fly [75], which significantly reduces computing costs compared to running separate simulations for each scale setting.
@@
The second source of uncertainty relates to the nonperturbative modeling of parton fragmentation into hadrons. These uncertainties arise from phenomenological models rather than first-principle calculations, leading to a significant model dependence. To evaluate fragmentation corrections, we use S HERPA ’s internal cluster hadronization model [63] and the Lund String model via an interface to P YTHIA 8 [64]. The symmetrized difference between these two models is taken as the uncertainty§§§Note this procedure can in principle result in uncertainties covering negative (differential) cross sections. These are clearly unphysical and we cut off the envelope at zero., using the cluster model as the central value.
@@
Another aspect of nonperturbative modeling involves tuning model parameters to data. We use the replica tunes from Ref. [76] to estimate uncertainties related to the tuning of final-state fragmentation parameters and the variations from Ref. [77] for parameters governing cluster decays involving beam remnants. This introduces two sets of variations within the cluster model itself, for which we take the respective envelopes for each histogram as the related uncertainty. As a result, three sources of nonperturbative uncertainty arise: model choice, final-state fragmentation tune, and beam fragmentation tune. Arguably the different sources of uncertainty should be combined. We, however, find that usually only one, in most cases the model choice, dominates the overall uncertainty. We hence only show the three sources independently, without a combination.
@@
Since estimating nonperturbative uncertainties requires independent Monte Carlo runs for each parameter set, and for both models, performing this study for the full MEPS@NLO, MEPS@LO and MC@NLO setups is infeasible. Instead, we derive the relative uncertainty based on parton showered leading-order matrix elements merged together in the MEPS@LO approach. A merged approach is still beneficial to ensure sufficient phase-space coverage. In the following we only treat the results of two tunes or models as different if we can statistically distinguish them at a level of 3 standard deviations, and otherwise indicate the statistical error that serves as an upper bound on the uncertainty.
@@
We make no attempt to directly apply this uncertainty to higher- or lower-precision simulations, though one could envision a procedure based on transfer matrices introduced in Ref. [78]. Additionally, we do not combine nonperturbative and perturbative uncertainty estimates but rather focus on comparing their relative sizes in the following analysis.
@@ 4 Event generation for HERA and MEPS@NLO in charged current DIS
Figure 2: Differential distributions for $Q^{2}$ and $x$ at LO, MC@NLO, MEPS@LO and MEPS@NLO in inclusive CC DIS at the HERA. The lower panels show the ratio between data and the merged MEPS@LO and MEPS@NLO predictions, while we omit the ratio for the LO and MC@NLO simulations for clarity. The full (hatched) band corresponds to variations of $\mu_{\text{F}},\mu_{\text{R}}$ ($Q_{\mathrm{cut}}$). Experimental data points are from Ref. [79].
The S HERPA framework has been extensively used in the past to describe neutral current DIS at HERA, see for example Ref. [12]. Early versions of S HERPA 3 were also used in a variety of recent event shape and jet analyses by the H1 Collaboration[80, 81, 82, 53, 83], usually finding satisfactory agreement with the data. A MEPS@NLO setup similar to the one used here was compared to resummed calculations for event shapes in Refs. [77, 84]. The latter also determined part of the current tune of nonperturbative parameters. With this, we consider the framework and MEPS@NLO to be sufficiently validated in neutral current DIS, and we focus in the following on the first detailed comparison to CC data from HERA.
We compare our Monte Carlo setups against multijet cross sections in CC DIS measured by the ZEUS Collaboration [79]. We consider beams of electrons at $27.5\,\text{GeV}$ and protons at $920\,\text{GeV}$. To model the proton contents, we use the NNPDF30_nlo_as_0118 set as discussed in Sec. 3.1.
In principle, a caveat for CC DIS is that, in contrast to NC DIS, it is impossible to use the final-state neutrino to reconstruct $q$. Therefore, the measurements of Eqs. (2) involve the reconstruction of the full hadronic final state under the condition of large missing $p_{t}$ from the invisible high-energy neutrino. The details of this procedure are discussed in Refs. [79, 85]. Since the data we compare to have been corrected for detector effects, we can ignore this complication and consider the recoil neutrino as known for the determination of $Q^{2}$, $x$ and $y$. In the analysis, events are selected following the experimental phase space cuts:
@@
Figure 3: Differential distributions for $\eta_{\mathrm{jet}}$ and $E_{T,\mathrm{jet}}$ at LO, MC@NLO, MEPS@LO and MEPS@NLO in inclusive CC DIS at the HERA. The lower panels show the ratio between data and the merged MEPS@LO and MEPS@NLO predictions, while we omit the ratio for the LO and MC@NLO simulations for clarity. The full (hatched) band corresponds to variations of $\mu_{\text{F}},\mu_{\text{R}}$ ($Q_{\mathrm{cut}}$). Experimental data points are from Ref. [79].
@@
Figure 4: Multijet distributions differential in the mass of the two (left) or three (right) leading jets at LO, MC@NLO, MEPS@LO and MEPS@NLO in inclusive CC DIS at the HERA. The lower panels show the ratio between data and the merged MEPS@LO and MEPS@NLO predictions, while we omit the ratio for the LO and MC@NLO simulations for clarity. The full (hatched) band corresponds to variations of $\mu_{\text{F}},\mu_{\text{R}}$ ($Q_{\mathrm{cut}}$). Experimental data points are from Ref. [79].
In Figs. 2 and 3, we reproduce the differential cross sections for inclusive-jet production as functions of $Q^{2}$, $x$, and leading-jet $\eta$ and $E_{\perp}$. Due to the relatively large $Q^{2}$ requirement, higher-order corrections are rather small, and we find the simulations at all accuracy levels to be consistent with each other and with the data. In particular, the lack of a sizable shift between the LO and MEPS@NLO predictions can be attributed to the high $Q^{2}$ cut. This requirement restricts the phase space to regions where merging effects remain small. In the lower panels, we show the ratio of the merged simulations to data. Since the central values for all Monte Carlo accuracies are very consistent, we do not show ratios for the LO and MC@NLO runs to increase legibility. For the two remaining runs, we also plot the band obtained from scale variations. Notably, while the central value is almost unchanged when going from MEPS@LO to MEPS@NLO, the scale dependence is reduced dramatically. This leads to some visible systematic trend of the MEPS@NLO simulation against data, especially in the transverse energy distribution, but the data are not precise enough there for a definite conclusion. Finally, we include an uncertainty band corresponding to the variations of the merging scale. In order to keep the plots legible, we only show the band for the MEPS@NLO calculation. The size is typically a few percent, though at the level of uncertainty indicated by the scale variations it can be the dominant effect for some bins. In Fig. 4, the dijet and three-jet invariant mass cross sections are presented. The different Monte Carlo samples are again very consistent with each other. For the dijet mass, some deviations appear for very small and very large masses. However only the deviations at very small masses have any statistical significance, and are driven by hadronization effects as we have confirmed by comparing parton and hadron level in our simulations. In the range $20\,\text{GeV}<M_{jj}<120\,\text{GeV}$, the Monte Carlo predictions are consistent with the data. In the three-jet mass case, prediction and data are consistent over the full measured range, albeit admittedly with significantly larger uncertainties on the data. The scale variations are reduced in both cases, although they remain sizable in the three-jet case. The merging scale uncertainties are of similar relative size as in the inclusive jet case, and hence much smaller than the scale variations even for the MEPS@NLO calculation.
@@ 5 Event generation for the EIC and its systematic uncertainties
Furthermore, we compare jet properties obtained using the $k_{t}$, anti-$k_{t}$, and C ENTAURO clustering algorithms, highlighting differences in jet multiplicities and leading jet transverse momentum. Finally, to probe the hadronic final-state, we study the 1-jettiness distribution, presented in slices of $Q^{2}$ and $x$.
@@
For both neutral and charged current DIS, we assume energies of $18$ and $275\,\text{GeV}$ for the lepton and proton beams respectively; this corresponds to the high-luminosity configuration of the EIC. We do not consider any polarization of the beams.
To arrive at realistic predictions, we discard particles outside an assumed detector acceptance of $\left|\eta\right|>4$ during the analysis and apply a cut on the inelasticity, keeping the interval $0.2<y<0.9$. Jets are clustered with a minimal transverse momentum of $5\,\text{GeV}$ and a jet radius of $R=1$. While we follow the setup of tradition studies like Ref. [16], the C ENTAURO algorithm presents new advantageous options that should be explored in dedicated studies [86].
@@ 5.1 Neutral current
Figure 5: Differential distributions for $Q^{2}$ (top left), $x$ (top right), $y$ (bottom left) and lepton transverse momentum $p_{\text{T}}^{\mathrm{lep}}$ (bottom right) in NC DIS at the EIC, comparing LO+PS, MC@NLO, MEPS@LO and MEPS@NLO predictions. The main panels show results with scale variation uncertainties for all but the LO+PS calculation. The first ratio panel displays the relative deviation from MEPS@NLO to highlight merging effects. The second ratio shows the relative scale uncertainties. The hatched band represents merging scale variation $Q_{\mathrm{cut}}$ in the MEPS@NLO sample.
We consider the neutral current setup first. In this case, we place a small virtuality cut, $Q^{2}>6\,\text{GeV}^{2}$. We present the global DIS observables $Q^{2}$, $x$, $y$, and $p_{\text{T}}^{\mathrm{lep}}$ in Fig. 5. For $Q^{2}$ and $x$, the merged calculations exhibit distinct behavior from the matched MC@NLO approach. In the $Q^{2}$ distribution (top left), we observe that for $Q^{2}<20\,\text{GeV}^{2}$, the MEPS@NLO prediction exceeds LO and MC@NLO of up to 30%, while showing a more modest 10% enhancement over MEPS@LO. This enhancement exceeds the scale uncertainty bands between MEPS@NLO and MC@NLO but remains within the MEPS@LO uncertainties. The predictions converge in the intermediate region $20\,\text{GeV}^{2}\leq Q^{2}\leq 5\cdot 10^{3}\,\text{GeV}^{2}$. At higher virtualities ($Q^{2}>5\cdot 10^{3}\,\text{GeV}^{2}$), we observe a systematic 10% deficit in the LO and MEPS@LO predictions relative to the NLO-accurate calculations, showing the importance of the virtual $\mathcal{O}(\alpha_{\text{s}})$ corrections in this regime.
Similar patterns can be seen in the $x$ and $p_{\text{T}}^{\mathrm{lep}}$ distributions (top and bottom right). Multijet merging effects dominate at $x<1.5\cdot 10^{-3}$ ($p_{\text{T}}^{\mathrm{lep}}<5\,\text{GeV}$), while NLO matching becomes crucial at $x>0.7$ ($p_{\text{T}}^{\mathrm{lep}}>65\,\text{GeV}$). The MEPS@NLO calculation shows reduced scale uncertainties compared to MEPS@LO in the low-$x$ and low-$p_{\text{T}}^{\mathrm{lep}}$ regions. It also shows smaller uncertainties than MC@NLO in the intermediate ranges of around $Q^{2}\approx 10^{2}\,\text{GeV}^{2}$ and $x\approx 10^{-2}$ as a consequence of calculating the $2\to 3$ process to NLO. The inelasticity distribution (bottom left) reveals good agreement between MEPS@NLO and MEPS@LO across the full range, while LO and MC@NLO predictions are uniformly suppressed by approximately 25%, though this deviation remains within the combined scale uncertainty bands.
The merging cut variations naturally are largest in the regions where merging has the largest impact, i.e. for small $Q^{2}$ and $x$, indicating a significant uncertainty on the way the DIS regime is effectively interpolated to the photoproduction region. The inelasticity distribution is affected everywhere by a flat shift from the scale variations. This is simply due to the fact that we integrate over the full $Q^{2}$ range, if we were to cut more aggressively to ensure a measurement cleanly in the DIS regime, we would expect a much smaller uncertainty.
@@
Figure 6: Jet multiplicity $N_{\mathrm{jet}}$ (top row) and leading jet transverse momentum $p_{T}^{\mathrm{jet}}$ (bottom row) distributions for anti-$k_{t}$ (left), $k_{t}$ (middle) and C ENTAURO (right) jet algorithms in NC DIS at the EIC. The bottom ratios in each row indicate, from top to bottom, the ratio of each prediction to the MEPS@NLO one, the uncertainty from scale variations and the uncertainty from nonperturbative model and tuning variations. The hatched band represents merging scale variation $Q_{\mathrm{cut}}$ in the MEPS@NLO sample. The error bars in the lowest panel indicate the statistical uncertainty of the difference between cluster and string model.
@@
Figure 6 presents the jet multiplicity and leading jet transverse momentum distributions for three different clustering algorithms. While MEPS@NLO and MEPS@LO predictions show good concordance, both MC@NLO and LO calculations yield systematically lower rates. Despite reduced scale uncertainties in MEPS@NLO compared to MEPS@LO, both merged calculations exhibit significant scale variations, reflecting the LO accuracy of higher-multiplicity configurations that dominate the low-$Q^{2}$ regime and that contribute significantly in these observables. Merging cut variations play a subleading role for these observables. It is also worth noting that the C ENTAURO clustering algorithm leads to smaller higher-order corrections in this analysis. Indeed, the LO simulation is unable to fill the whole phase space. For the $k_{t}$ and anti-$k_{t}$, the MC@NLO sample is still off, compared to the merged ones, by a $K$-factor of about 1.5 for the low-$p_{T}$ region. Conversely, the C ENTAURO algorithm shows much smaller $K$-factors, in fact only a factor of 1.2 for the low-$p_{T}$ region with respect to the MEPS@NLO prediction.
@@
Hadronization uncertainties are generally smaller than the perturbative uncertainties and, as one might have expected, are largest for small $p_{T}$ (up to $5\%$) and for large jet multiplicities $(1-2\%)$. In the latter, we find that only a variation of the fragmentation model induces a significant effect, while replica tunes within the cluster model are negligible. Instead, in the $p_{T}$ spectrum, fragmentation model uncertainty is still the largest nonperturbative contribution for the anti-$k_{t}$ and $k_{t}$ jets, while variations from replica tunes are larger than model ones for C ENTAURO jets, although both are especially small. Beyond small values of $p_{T}$, the difference between hadronization models becomes unresolvable within the accumulated statistics. Instead, the combined statistical uncertainty from the two models is taken as an overestimate. These values are still comparable or subdominant to the perturbative uncertainties in the same part of the spectrum.
@@
Figure 7: Differential distributions of 1-jettiness $\tau$ in different bins of $Q^{2}$ and $x$ in NC DIS at the EIC. The total MEPS@NLO cross section for each $x$-$Q^{2}$ bin is reported on top of the corresponding plot.
@@
We finally consider 1-jettiness, an event shape variable that is sensitive to the underlying parton dynamics and can be used, e.g., for global fits of the strong coupling constant $\alpha_{\text{s}}$ [87]. In Ref. [88], predictions of 1-jettiness at the EIC were computed with NNLO+N3LL accuracy. Here, we present in Fig. 7 an overview of the 1-jettiness $\tau$ across bins of $x$ (left to right) and $Q^{2}$ (top to bottom). For the sake of presentation we omit the ratio panels here, and the full plots can be found in the Appendix. Overall, the shape of the distributions agrees well between the different configurations. Similar to the behavior at the recent H1 measurement [82], we observe that the high-$Q^{2}$ has a pronounced peak at small $\tau$ values, while the low-$Q^{2}$ distributions tend to higher values, which underlines the importance of high-multiplicity matrix elements in the latter regime.
@@
For a more thorough discussion, we show the full breakdown of ratios and uncertainties in Fig. 8 for three different $x$ regions in the $10<Q^{2}/\,\text{GeV}^{2}<31.6$ range. Plots for the other $x$-$Q^{2}$ ranges in Fig. 7 can be found in Appendix A.1. We observe that the nonperturbative uncertainties are non-negligible for small $\tau$ values and of similar size as the MC@NLO and MEPS@NLO scale uncertainties in that region. The same qualitative behavior is visible in both the tuning variations and the difference between string and cluster models, while the latter are significantly larger. This is the case for both tuning uncertainties and variations of the hadronization model, where the latter are dominating. For larger values of 1-jettiness, either of the considered nonperturbative uncertainties are negligible, consistent with the observations in Ref. [77]. Furthermore, the nonperturbative uncertainties are—unsurprisingly—largest for small values of $Q^{2}$ and $x$. The shape of the distributions at small $Q^{2}$ underlines the importance of multiparton topologies in this region. We find that variations in $Q_{\text{cut}}$ primarily affect the overall cross section and not the distribution itself, apart from regions where NP effect are completely dominant. Thus, we choose to omit them and show the 1-jettiness distributions in each $Q^{2}$$x$ bin with no normalizing factor to emphasize the effect of the different MC setups.